OpenStudy (ashy98):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 A.) -2y + 6 B.) 2y - 6 C.) 2y + 6 D.) -2y - 6

OpenStudy (ashy98):

@Awolflover1 @Angelus

Awolflover1 (awolflover1):

What do you think ?

Awolflover1 (awolflover1):

-x+2y=-6 -x=-2y-6 x=2y+6

OpenStudy (ashy98):

c

Awolflover1 (awolflover1):

Correct :)

OpenStudy (ashy98):

thank you !!!(: